OpenStudy (omw2fyb):

1. Kevin ran 4 miles more than Steve ran. The sum of their distances is 26 miles. How far did Steve run? The domain of the solution is {0, 4, 11, 22}. a. Define your variable. b. Write algebraic expressions for the distances each person ran. c. Write an equation that can help you solve the problem. d. Solve the equation. e. Check your answer.

likeabossssssss (likeabossssssss):

\(\Huge\color{#EB00FF}{\text{WELCOME}}\) \(\Huge\color{blue}{\text{TO}}\) \(\Huge\color{green}{\text{OPEN}}\)\(\Huge\color{purple }{\text{ STUDY!!!!!!!!!!!}}\) \(\Huge\heartsuit\)

OpenStudy (festinger):

S=x K=x+4 S+K=26=2x+4 Solving 2x+4=26 will get you x, which is the distance steve covered.